OpenStudy (anonymous):

write the simplest polynomial function with the zeros 2-i,√5, and -2.

OpenStudy (alexandervonhumboldt2):

hint: (x-(2-isqrt5))*(x-(-2))=?

OpenStudy (solomonzelman):

In general, the simplest polynomial function with zeros, \([\rm part~1]\) \(\color{#000000 }{ \displaystyle a }\) \([\rm part~2]\) \(\color{#000000 }{ \displaystyle a,~b }\) \([\rm part~3]\) \(\color{#000000 }{ \displaystyle a,~b,~{\tt \small and}~c }\) \(\rm And~so~forth{\bf~.....}\) is going to be: \([\rm part~1]\) \(\color{#000000 }{ \displaystyle f(x)=x-a }\) \([\rm part~2]\) \(\color{#000000 }{ \displaystyle f(x)=(x-a)(x-b) }\) \([\rm part~3]\) \(\color{#000000 }{ \displaystyle f(x)=(x-a)(x-b)(x-c) }\) Reasoning/because \([\rm part~1]\) \(\color{#000000 }{ \displaystyle f(x)=0 }\) when evaluated at \(\color{#000000 }{ \displaystyle x=a }\) \([\rm part~2]\) \(\color{#000000 }{ \displaystyle f(x)=0 }\) when evaluated at \(\color{#000000 }{ \displaystyle x=a,~{\small \tt or}~b }\) \([\rm part~3]\) \(\color{#000000 }{ \displaystyle f(x)=0 }\) when evaluated at \(\color{#000000 }{ \displaystyle x=a,~b,~{\small \tt or}~c }\)

OpenStudy (solomonzelman):

If something is not making sense, then ask...

OpenStudy (solomonzelman):

NOTE: This applies to absolutely all numbers \(a, b, c ~....\) (Complex \(\mathbb{C}\), or Real \(\mathbb{R}\))

OpenStudy (anonymous):

it makes sense, thank you @SolomonZelman and @AlexandervonHumboldt2

OpenStudy (solomonzelman):

No problem .

OpenStudy (mathmale):

Principle to follow here: If a is a root of a given polynomial, then (x-a) is a factor. Look at the given roots and turn them into factors. Note that if you have a complex root, you must find the conjugate of that root; both must be turned into factors. Show your work, please.

OpenStudy (mathmale):

If one root is -2, what is the corresponding factor?

OpenStudy (solomonzelman):

Mathmale, there is only one complex root. (It's possible for a polynomial to have complex coefficients)